lazyking4life:

Find the measure of x. (to the nearest tenth) A) 25.7 B) 29.6 C) 38.1 D) 68.2

lazyking4life:

1 attachment
snowflake0531:

You have the adjacent and opposite side, which trig ratio do you use \[\sin(\theta) = \frac{opposite}{hypotenuse}\] \[\cos(\theta) = \frac{adjacent}{hypotenuse}\] \[\tan(\theta) = \frac{opposite}{adjacent}\]

lazyking4life:

huh

lazyking4life:

wym

@snowflake0531 wrote:
You have the adjacent and opposite side, which trig ratio do you use \[\sin(\theta) = \frac{opposite}{hypotenuse}\] \[\cos(\theta) = \frac{adjacent}{hypotenuse}\] \[\tan(\theta) = \frac{opposite}{adjacent}\]
wym

snowflake0531:

Theta is the angle, opposite is the opposite side, adjacent is the side that is not the hypotenuse, and hypotenuse is hypotenuse I meant earlier that we needed the adjacent side and hypotenuse And so with theta, adjacent and hypotenuse, we know that we have to use the cosine trig ratio What grade are you in?

snowflake0531:

Credits to an anonymous person:

1 attachment
lazyking4life:

@snowflake0531 wrote:
Theta is the angle, opposite is the opposite side, adjacent is the side that is not the hypotenuse, and hypotenuse is hypotenuse I meant earlier that we needed the adjacent side and hypotenuse And so with theta, adjacent and hypotenuse, we know that we have to use the cosine trig ratio What grade are you in?
10th